OpenStudy (anonymous):

Solve the proportion 5/x = 4/8 A. 2 B. 6 C. 8 D. 10

OpenStudy (cmtboy2):

8

OpenStudy (swishking19):

No, it's D.

OpenStudy (anonymous):

no thats not the awser

OpenStudy (anonymous):

5x=48 Step 1: Cross-multiply. 5x=48 (5)*(8)=4*x 40=4x Step 2: Flip the equation. 4x=40 Step 3: Divide both sides by 4. 4x4=404 x=10 Answer: x=10

OpenStudy (cmtboy2):

the answer is c.8

OpenStudy (swishking19):

No. 4/8 is one half, so it would have to be 5/10.

OpenStudy (swishking19):

Because 5/10 is also one half.

OpenStudy (ddos_dragon):

NO Direct answers.

OpenStudy (anonymous):

i solved it as u can see

OpenStudy (cmtboy2):

ok you are right

OpenStudy (ddos_dragon):

You didn't walk @tjl3k Through it.

OpenStudy (swishking19):

Yeah, you're right.

OpenStudy (ddos_dragon):

Next time be advised not to break the rules you agreed to follow.

OpenStudy (swishking19):

Thanks. I am new here, so yeah.

OpenStudy (ddos_dragon):

:) Make sure and Read the CoC as well as the TaC.

OpenStudy (swishking19):

Okay I will.

OpenStudy (ddos_dragon):

They are short, and will keep you safe.

OpenStudy (ddos_dragon):

From.. suspension, that is

OpenStudy (swishking19):

Lol yeah I guess.